State whether the following statements are true or false. Justify yourself with the help of a short proof or a counter example.
i) y′ + P(x) y = Q(x) y^n is a linear equation for integral values of n.

ii) y = 0, is a singular solution of the differential equation 27y-8(dy/dx)^3=0

iii) Equation x^2 ( y − px) = yp^2 is reducible to clairaut’s form
1
Expert's answer
2020-04-03T13:24:46-0400

State whether the following statements are true or false. Justify yourself with the help of a short proof or a counter example.

i) y′ + P(x) y = Q(x) y^n is a linear equation for integral values of n. 

SOLUTION:-

"y' + p\left( x \right)y = q\left( x \right){y^n}\quad (*)"

where p(x) and q(x) are continuous functions on

the interval we’re working on and n is a real number.

divide the differential equation by "y^ \nn" to get

"{y^{ - n}}\,y' + p\left( x \right){y^{1 - n}} = q\left( x \right)"

We are now going to use the substitution

"v = =y ^{1-n}"

to convert this into a differential equation in terms of v

As we’ll see this will lead to a differential equation that we can solve. So

"v' = \left( {1 - n} \right){y^{ - n}}"

Now, plugging this as well as our substitution into the differential equation gives,

"\frac{1}{{1 - n}}v' + p\left( x \right)v = q\left( x \right)"

This is a linear differential equation that we can solve for v

and once we have this in hand we can also get the solution to the

original differential equation by plugging v.

 back into our substitution and solving for y

So the equation is reduced to a linear equation.

If n=0 equation (*) is a linear equation, in our case

n=2 equation is not linear equation, but reduced to a linear equation.


Look at an example.

"y' + \frac{4}{x}y = {x^3}y^2"

So, the first thing that we need to do is get this into

the “proper” form and that means dividing everything by "y ^\n2" .. Doing this gives

"{y^{ - 2}}\,y' + \frac{4}{x}{y^{ - 1}} = {x^3}"


The substitution and derivative that we’ll need here is

"v = {y^{ - 1}}\hspace{0.25in}v' = - {y^{ - 2}}y'"

With this substitution the differential equation becomes

"\u2212v' \n\n + {4 \\above{2pt} x}\n\n\u200b\t\n v=x ^\n3"

Here’s the solution to this differential equation

"v' - \frac{4}{x}v = - {x^3}\hspace{0.25in}\,\,\,\, \Rightarrow \hspace{0.25in}\\\mu \left( x \right) = {{\bf{e}}^{\int{{ - \,\,\frac{4}{x}\,dx}}}} = {{\bf{e}}^{ - 4\,\,\ln \left| x \right|}} = {x^{ - 4}}"

"\u222b(x^{ \u22124} v) 'dx=\u2212ln\u2223x\u2223+c\u21d2\nv(x)=cx^ 4\u2212x ^4lnx"

So, to get the solution in terms of y

 all we need to

do is plug the substitution back in. Doing this gives,

"y ^{
−1}
 =x ^
4
 (c−\ln{x})"

ii) y = 0, is a singular solution of the differential equation 27y-8(dy/dx)^3=0 

General solution to the equation is: "y=(x+C)^\frac{3}{2}" .

The solution y=0 cannot be obtained from the general solution for any value of constant with C,

hence y=0  is a singular solution of this equation.

iii) Equation x^2 ( y − px) = yp^2 is reducible to clairaut’s form

Let's put "X=x^2;\quad Y=y^2" then "\u200b" "p={x\above{2pt} y}{dY \above{2pt} dX}" .. Let's denote "P= {dY\above{2pt} dX}"


Then the equation can be rewritten in form

"X(y−x^ 
2{P\above{2pt} y}=y{x^2\above{2pt} y^2}P^2"

By multiplying both sides with y, we assume

"X(y^ \n2\n \u2212x ^\n2\n P)=x^ \n2\n P ^\n2"

or

"X(Y\u2212XP)=XP ^\n2"

Therefore

"Y=XP+P \n^2"

which is now in Clairaut’s form

The solution got by just replacing P by constant c.

Hence

"Y=cX+c^2"

"or \,

y^2=cx^2+c^2"